OpenStudy (anonymous):

Simplify 5(a - b) + 6(a + b) - 7(a - 2b)

OpenStudy (khaakenajaf):

=5a -5b+6a+6b-7a+14b =5a+6a-7a+6b+14b-5b =11a-7a+20b-5b =4a+15b

OpenStudy (anonymous):

thankz